You are here

Project Updates for collection: all

Search

Search tips:

  • This form will search for words in the title OR the description. If you would like to search for the same term(s) across both the title and description, enter the same search term(s) in both fields.
  • This form will search for any of the words you enter in a field, not the exact phrase you enter. If you would like to search for an exact phrase, put double quotes (") around the phrase. For example, if you search for Bike Path you will get results containing either the word Bike OR the word Path, but if you search for "Bike Path" you will get results containing the exact phrase Bike Path.
  1. Fiber-optic cable intact

    Hi everyone,

    I can report today that the fiber-optic cable is intact and the test returned a signal along the entire 235 m cable. What you see in the photograph is data along the cable going down and coming back up (470 m total).

    We have ordered the drop box to store the fiber-optic cable, and a hole has been dug for it. The box will be installed by F&S after the holidays. The fencing is being left around the site until this is done.

    Wishing you all the best this holiday season.

    Andy

    Attached Files: 
  2. Borehole sealed

    Hi Everyone, We successfully installed the geothermal loop and fiber optic cable today. The borehole is now grouted closed. My appreciation goes out to the group of colleagues and staff from the ISGS that assisted the drill crew to insert the 400-feet of tubing in the ground.

    The drill crew will be back in the morning to clean up the site and move equipment offsite. We still need to install a drop box to store the fiber-optic cable and access the geothermal loop. The box has been ordered and if it arrives in the morning, we may be able to get it installed before the holidays. In about a month we will start measuring the ground temperature and run a DTRT test once the grout cures.

    Happy Holidays!!

    Andrew

  3. Installing fiber optics

    Hi everyone, We had a successful day of collecting geophysical logs in the borehole. I have seen print outs of this excellent data, and is an important part of the geologic characterization. Thanks goes out to Tim Young for coming out of retirement to help us complete the work.

    Tomorrow we will install the U-bend and fiber-optic cable in the borehole. Our task is to sink a 400-foot log tube filled with water (weighing ~600 lbs) down the hole. For those nearby you may want to drop by for a visit. Once this is done, we will proceed to grout the borehole closed.

    Cheers,

    Andy

  4. Last drilling update

    Hi everyone,

    We finished drilling today and advanced the borehole to a depth of 385 feet.

    Tomorrow we will collect geophysical logs of the borehole. If you want to drop by we should be in full swing by 11 am. It will take some time to clean out any material that has collapsed in the borehole overnight and make sure the hole is full of mud before logging the open hole.

    Best,

    Andy

  5. Weekly update

    Associated Project(s): 

    All, Last week was slow, per usual this time of year. I had a training session for two new staffers. They’ll start in the new year as they have finals this week.

    Bike count is hovering around 15-20 with a few safety checks still in the queue. A couple staffers expressed interest in morning shifts this week so if that happens, hopefully it’ll result in a few more builds being ready before winter break.

    This week I will continue work on the Student Staff Manual and will line up some new volunteer tasks since the regular items are getting ticked off the list.

    The numbers:

    Visitors: 51
    Sales: $117.50
    tire/tubes: 6 for $23


    Thanks!

     

    -        Jake Benjamin
    Campus Bike Center Manager

  6. Weekly Updates for Zero Waste

    Associated Project(s): 

    Hi Pete and Shawn--

    This past week, I posted Morgan's message about the change to the glove recycling program to the iCAP portal. I also followed up on a response from the Kimberly-Clark representative. Morgan has provided additional clarification about the changes and accepted her invitation for further dialog.

    Best regards,

    Marya Ryan 

  7. 18F Semesterly Report - BIF Rooftop Solar

     

    The BIF project was initially scheduled to begin in January 2018, but was on hold until F&S and the College received word from the state regarding the general contractor award.  The kickoff meeting took place in March 2018 and a planning meeting followed in April. 

     

    On Monday, May 21st, construction kicked off and is expected to be complete summer 2019.  Thus far, the following work has been completed (64% complete):  removal of rooftop garden; insallation of steel, metal, and concrete decking; roof membrane installed; installation of roof drain lines; installation of exterior and interior walls;  window and ductwork installation; HVAC updates/upgrades; above ceiling mechanicals and sprinkler lines installed; GWB taping and sanding with painting underway. 

     

     

  8. Glove Recycling - Program Changes

    Associated Project(s): 

    Morgan White sent the following message to labs participating in the glove recycling program on Dec. 13, 2018:

    Hello Glove Recycling Contacts,

     

    Thank you for your support of the Illinois Climate Action Plan (iCAP) waste reduction efforts, in connection with the Kimberly-Clark Professionals RightCycle program.  We are very appreciative of the initiative and time you all have dedicated to glove recycling to help reduce landfill waste from the University of Illinois over the past three years. I’m also emailing the people who have expressed interest in joining the program, so you are aware of these changes.

     

    I am writing to let you know that we will halt glove recycling for research labs and teaching labs on campus, as of the end of this calendar year.  The RightCycle program is only intended for recycling non-hazardous materials, and the majority of our labs are using gloves as protection against hazards or potential hazards.  Therefore, we will continue to participate in RightCycle only with gloves used for food service, starting in 2019.  If you are collecting gloves to recycle from a lab on campus, please do not drop off gloves and do not ship them to Kimberly-Clark after December 31, 2018.

     

    If you and your colleagues in research and teaching labs would like to continue actively supporting campus sustainability, please join the Certified Green Lab Program at https://go.illinois.edu/greenlab, coordinated through the Institute for Sustainability, Energy, and Environment (iSEE).

     

    Please let me know if you have any questions about glove recycling, and have a happy holiday season.

     

    Thanks,

    Morgan

  9. Weekly Updates for Zero Waste

    Associated Project(s): 

    The updates below apply to the week ending Dec. 9, 2018.

    Hi Pete and Shawn--

    This past week, zero-waste activity focused on a major change to the glove recycling program. Morgan and I spoke with Monica Miller on Friday and heard concerns about the potential for labs (especially research labs) to inadvertently place hazardous materials in the recycling stream. Monica pointed out the complexities of compliance, especially when staff turns over at a lab. The determination was to stop glove recycling at labs. There may be some exceptions that DRS would need to determine on a case-by-case basis.

    Morgan is drafting a message to the labs to tell them of the changes, which will take effect Dec. 31. Per Monica's request, I have contacted Serenity Desmond at Noyes and the Chem Annex to ask her to contact Monica about the change. Monica is familiar with ISTC's practices and is comfortable with their continued participation. There is no issue with Dining Services or food venues continuing to participate.

    I have contacted the BER students who were trying to find the names of teaching coordinators so that DRS could talk with them under the previous plan. I told them to send the information they have already collected but that they can stop. Morgan and I hope to enlist them to help expand the program in food venues, and possibly to promote RightCycle to restaurants in the community.

    Best regards,
    Marya Ryan

  10. Daily update

    Hi everyone, As expected our rate of drilling today slowed when drilling the shale bedrock. The crew was able to drill an additional 25 feet to a depth of 355 feet.

    Next week, since Monday will be the last day we have to drill the hole deeper, we have decided to stop coring and just drill straight through so that we can reach the target depth of 450 feet. Tuesday is needed to clean out the borehole so that geophysical logging can be done on Wednesday. On Thursday and Friday, we will install the geothermal loop and fiber optic cable and grout the hole closed.

    Have a great weekend.

    Andy

  11. Daily update

    Hi Everyone, We are starting out Wednesday morning at 294 feet, about 5 feet into the bedrock. The crew will be widening the hole this morning from 5 inches to 5.9 inches before going deeper into bedrock. Looks like a nice day for drilling!!

    Andy

  12. Daily update

    Hi everyone,

    After a morning of thawing out the drilling system from well below freezing temperatures on the weekend, we were able to reach 240 feet. With the top bedrock being <60 feet, it is likely they will reach it on Tuesday.

    My colleague Jason Thomason will be at the site on Tuesday while I am out of town.

    Best,

    Andy

  13. Weekly update

    All, Visitor numbers are down, sales are down, builds are ever-creeping up: It must be winter!

    We’re only averaging a few actual work-on-my-bike visitors per day now, and I imagine it’ll get even slower the closer we get to winter break. The student staff has responded positively to the shift in shop build protocol away from piecemeal jobs and towards being wholly responsible for the bike, start to finish. I don’t think it has actually positively affected the quality of their work—yet—but they seemed more engaged and invested in the task at hand. 

    Not sure the numbers on it but I feel like I’ve had a pretty steady flow of volunteers in for either community service or volunteer memberships. Not too many volunteer duds, as most seem pretty jazzed and impressed with the space/DIY idea.

    This week I’ve got a meeting with the GM of VeoRide and separately, a TBP-UIUC meeting, as well as student staff training towards the end of the week.

    The numbers:

    Visitors: 43

    Sales: $144.50

    Memberships: 2 for $60

     

    Thanks!

     

    -        Jake Benjamin
    Campus Bike Center Manager

  14. Daily update

    Hi everyone,

    We had a successful day 2 after deicing the system. We are now down to 165 feet and heading into the oldest glacial strata. If you are following along with the log I sent, we are about 170 feet on it.

    We will start up again Monday morning.

    Best,

    Andy

  15. daily update

    Hi everyone,

     

    After day 1 we have drilled to a depth of 90 feet. This is excellent progress with snow much of the day, and not getting in the ground until late morning.

    So far, the geology is similar to previous drilling in the area (see log attached).

     

    Cheers,

    Andy

Pages